# ChangeMamba: Remote Sensing Change Detection with Spatio-Temporal State Space Model

### v1

@@ IV-B1 Implementation Details
All three of our proposed architectures are implemented in Pytorch. Depending on the size and depth of the encoder network, our proposed MambaBCD, MambaSCD, and MambaBDA architectures are available in Tiny, Small, and Base versions [30]. As shown in Table II, the difference between them is the number of VSS blocks inside each stage and the number of channels in the feature map. For datasets other than the SYSU dataset, the multi-temporal image pairs and associated labels are cropped to 256256 pixels for input to the network, and then the data with the original size is inferred using the trained networks on the test set. During training, we optimized the network using the AdamW optimizer [57] with a learning rate of 1 and weight decay of 5, without any learning rate decay strategy. The batch size is set to 16. Except for the SYSU dataset where we set the number of training iterations to 20000, we set the number of training iterations to 50000 on the remaining four datasets. The random rotation, left-right, and top-bottom flip are used as training data augmentation methods. Our source code will be open-sourced111https://github.com/ChenHongruixuan/MambaCD for community reproduction and subsequent research.
IV-B2 Assessment Metrics
For evaluating model performance across different CD subtasks, we employ a set of metrics tailored to the specific requirements of BCD, SCD, and BDA.
In assessing BCD, we use the following metrics:
Recall (Rec): The proportion of actual positives that were correctly identified is measured by the recall, defined as
(11) where represents true positives and denotes false negatives.
Precision (Pre): Reflecting the accuracy of positive predictions, precision is defined as
(12) with standing for false positives.
Overall Accuracy (OA): The ratio of correctly predicted observations to the total observations is represented by
(13) where indicates true negatives.
F1 Score (F1): Balancing both precision and recall, harmonic mean, given by
(14) -
(15) -
Cohen’s Kappa (): Assessing the agreement between two raters beyond chance, Cohen’s Kappa is given by
(16) where is the observed agreement, and is the expected agreement by chance.
For SCD, the metrics used include:
Overall Accuracy (OA) and F1 Score (F1): As defined for BCD.
(17) aiming to provide a balanced measure across different classes.
For BDA, we follow the protocol of xView2 Challenge to split BDA into two subtasks, i.e., building localization and damage classification tasks. The following metrics are adopted to measure the performance of each subtask and the overall BDA performance:
F1 Score for Localization () and Classification (): These metrics assess accuracy in building localization and damage classification, respectively.
Overall F1 Score (): This reflects the combined performance in both building localization and damage classification tasks.
F1 Score for Different Damage Level Classes (): This evaluates the model’s performance across various damage extents.
IV-C Comparison Methods
To rigorously assess the performance of Mamba in CD tasks, we strategically selected a diverse range of methods based on CNN and Transformer architectures for benchmarking across three distinct tasks: BCD, SCD, and BDA.
For BCD, our comparison includes a suite of CNN-based methodologies: FC-EF [16], FC-Siam-Diff [16], FC-Siam-Conc [16], and SiamCRNN [13] (utilizing ResNet architectures ranging from ResNet-18 to ResNet-101), alongside SNUNet [32], DSIFN [17], HANet [33], and CGNet [34]. On the Transformer front, we compare against ChangeFormer (versions 1 to 4) [22], BIT (with ResNet-18 to ResNet-101 as the backbone for feature extraction) [20], TransUNetCD [25], and SwinSUNet [24].
In the SCD arena, CNN-based competitors encompass HRSCD (variants S1 to S4) [35], ChangeMask [18], SSCD-1 [39], Bi-SRNet [39], and TED [45]. For Transformer-based methodologies, we evaluate against SMNet [44] and ScanNet [45].
Lastly, for the BDA task, we selected the xView2 baseline [42], Siamese-UNet (the winning solution of the xView2 challenge) [42], MaskRCNN [43], and ChangeOS [3] (employing ResNet-18 to ResNet-101) as CNN-based benchmarks. DamFormer [23] stands as the solitary Transformer-based comparison method.
For the above methods, if the detection performance was reported in the original paper on the corresponding dataset, we directly adopt the accuracy in the original paper. If not, we train and test the method on the corresponding dataset based on the hyperparameters recommended in the original paper, using the loss function and data augmentation methods consistent with our method.
IV-D Benchmark Comparison in Three CD Subtasks
@@ IV-D1 BCD
Tables III to V list the performance in BCD of the proposed MambaBCD architecture and some comparison methods on the three benchmark datasets. In these tables, we differentiate between the types of these BCD architectures into CNN-based methods, Tranformer-based methods, and our Mamba-based methods. It can be seen that our approaches significantly outperform either the CNN-based approaches or the Transformer-based architectures. The proposed MambaBCD-Base and MambaBCD-Small architectures achieve the highest and second-highest OA, F1, IoU, and KC for both class-agnostic CD (SYSU dataset) and single-object CD tasks (LEVIR-CD+ and WHU-CD datasets), fully demonstrating the potential of the Mamba architecture for the BCD task. Fig. 7 to Fig. 9 then show some binary change maps predicted by our three architectures on the test sets of the three datasets. It can be seen that the proposed methods accurately detect changes with varying types, scales, and numbers contained in these image pairs, very close to the change reference maps.
@@
| MamabaSCD-Base | 88.00 | 63.98 | 73.47 | 23.94 |
@@
| MamabaSCD-Tiny | 19.44 | 63.72 | |
| MamabaSCD-Small | 51.82 | 137.10 | |
| MamabaSCD-Base | 87.47 | 201.85 |
IV-D2 SCD
Table VII compares the SCD performance of the proposed MamabaSCD architectures and representative SCD approaches. Compared to the BCD task, the SCD task requires the change detector capable of effectively modeling not only the spatio-temporal relationships of multi-temporal images, but also learning representative and discriminative representations for accurate land-cover mapping. From Table VII, we can see that the MambaSCD architecture can satisfy both requirements well. The MambaSCD-Small architecture surpasses the Transformer-based SOTA approach, ScanNet, on all four metrics of the SCD task, with OA of 88.38, F1 of 64.10, mIoU of 73.61, and SeK of 24.04 . Note that our methods have achieved the current accuracy without employing loss functions such as those proposed for the SCD task in ScanNet [45]. Table VIII further lists the number of parameters and GFLOPs of these methods. Compared to the Transformer-based ScanNet method, the GFLOPs of our three architectures are lower. Compared to SMNet, another Transformer-based method, the proposed MamabaSCD-Tiny can achieve higher SeK values with lower number of parameters and GFLOPs. These results demonstrate the effectiveness of the proposed methods and the potential of the Mamba architecture on the SCD task.
@@ IV-D3 BDA
| xView2 Baseline | 80.47 | 3.42 | 26.54 | 66.31 | 14.35 | 0.94 | 46.57 | |
| Siamese-UNet | 85.92 | 65.58 | 71.68 | 86.74 | 50.02 | 64.43 | 71.68 | |
| MaskRCNN [43] | 83.60 | 70.02 | 74.10 | 90.60 | 49.30 | 72.20 | 83.70 | |
@@ IV-E Different Spatio-Temporal Fusion Methods
Modelling spatio-temporal relationships is significant to CD tasks [36, 13, 18]. To demonstrate the effectiveness of our proposed CST-Mamba module, we compare it with several commonly used spatial-temporal modeling methods in CD. They are concatenation operations with feature pyramid network (FPN) [18] (baseline), temporal-symmetric transformer based on 3D convolution layers (TST) [18], RNN-based method [13], and Tranformer-based method [20]. Table X lists the performance of these modeling methods on the three CD subtasks. On the simpler BCD tasks, Transformer does not yield significant advantages in performance compared to RNN. However, on the more difficult SCD and BDA tasks, the advantages of the Transformer architecture are illustrated. Compared with these current representative spatio-temporal relationship modeling approaches, our proposed cross-temporal modeling approach based on Mamba, which inherits the advantages of RNN and Transformer, shows superior performance on the three CD tasks of BCD, SCD, and BDA. Compared to the Transformer-based spatio-temporal modeling approach, our method improves the F1 value by 0.77 on BCD, SeK by 0.6 on the SCD task, and by 1.10 on the BDA task.
In addition, We further visualize the features extracted by the MambaBCD architecture in Fig. 12. In the features extracted by the encoder, the features in the region of high response values are not necessarily regions of interest, i.e., changed regions. Whereas, in the decoder, after the spatio-temporal relationships are gradually modeled by the proposed CST-Mamba modules, changed regions gradually exhibit high response values.
@@ IV-F Comparison to Other Backbone Networks
| Encoder | OA | F1 | IoU | KC | Params | GFLOPs |
| ResNet-101 | 91.10 | 81.10 | 68.22 | 75.28 | 46.90 | 131.36 |
| EfficientNet-B5 | 91.60 | 82.26 | 69.86 | 76.84 | 6.55 | 53.62 |
| MixFormer-v3 | 87.70 | 75.00 | 60.00 | 66.87 | 47.03 | 82.98 |
| Swin-Small | 92.09 | 82.22 | 69.80 | 77.16 | 36.22 | 81.72 |
| Mamba-Small | 92.35 | 82.83 | 70.70 | 77.94 | 49.94 | 114.82 |
@@
Finally, we simply compare the Mamba with some representative backbone networks in CD tasks on the SYSU dataset in Table XI. It can be seen that Mamba outperforms the two Transformer backbone networks, Swin-Transformer and MixFormer. This is because, in order to reduce the computational overhead of self-attention operation, these two Transformer backbone networks either adopt a local attention mechanism (Swin-Transformer) [26] or a strided convolution layer (MixFormer) [27] to reduce the size of the feature maps, which will negatively affect the learning of global contextual information. The Mamba architecture, on the other hand, can reduce its computational consumption to a linear relationship with the number of tokens, thus eliminating the need for these methods, which cause information loss, and allowing the fully learning of global context information. As for the remote sensing change detection task, due to the multi-scale phenomenon of land-cover features and the differences in spatial resolution of different sensors, the sizes of changed objects often vary greatly. Therefore, the features extracted by Mamba through fully learning the global contextual information, are particularly important for the subsequent detection of these changed objects with different sizes.

### v2

@@ IV-B1 Implementation Details
All three of our proposed architectures are implemented in Pytorch. Depending on the size and depth of the encoder network, our proposed MambaBCD, MambaSCD, and MambaBDA architectures are available in Tiny, Small, and Base versions [38]. As shown in Table II, the difference between them is the number of VSS blocks inside each stage and the number of channels in the feature map. For datasets other than the SYSU dataset, the multi-temporal image pairs and associated labels are cropped to 256256 pixels for input to the network, and then the data with the original size is inferred using the trained networks on the test set. During training, we optimized the network using the AdamW optimizer [66] with a learning rate of 1 and weight decay of 5. The batch size is set to 16. Except for the SYSU dataset, for which we set the number of training iterations to 20000, we set the number of training iterations to 50000 on the remaining four datasets. The random rotation, left-right, and top-bottom flip are used as training data augmentation methods. Our source code will be open-sourced111https://github.com/ChenHongruixuan/MambaCD for community reproduction and subsequent research.
@@
IV-B2 Evaluation Metrics
For evaluating model performance across different CD subtasks, we employ a set of metrics tailored to the specific requirements of BCD, SCD, and BDA. In evaluating the BCD performance of the model, we use the following metrics:
Recall (Rec): The proportion of all true positive pixels correctly identified as positive
(11) where represents true positives and denotes false negatives.
Precision (Pre): Reflecting the proportion of true positive pixels among all pixels identified as positive
(12) with standing for false positives.
Overall Accuracy (OA): The ratio of correctly predicted pixels to total pixels is expressed as
(13) where indicates true negatives.
F1 Score (F1): Balancing both precision and recall by calculating their harmonic mean, given by
(14) -
(15) -
Cohen’s Kappa (KC)): Assessing the agreement between two raters beyond chance, Cohen’s Kappa is given by
(16) where is the observed agreement, and is the expected agreement by chance. Their detailed definition can be found in [14]
For SCD, we follow the existing setup in the literature [55] to use the following metrics:
Overall Accuracy (OA) and F1 Score (F1): As defined for BCD but extended into multi-class cases.
(17) where is the number of semantic categories. mIoU can provide a balanced measure across different semantic categories.
For BDA, we follow the protocol of xView2 Challenge [51] to split BDA into two subtasks, i.e., building localization and damage classification tasks. The following metrics are adopted to measure the performance of each subtask and the overall BDA performance:
F1 Score for Building Localization (): This metric assesses model’s accuracy in building localization task.
F1 Score for Different Damage Level Classes (): This evaluates the model’s performance across various damage levels.
F1 Score for Damage Classification (): This metric evaluates the overall performance of the model on the damage classification task and is defined as the harmonic mean of
(18) where the value of in the xBD dataset is 4, from 1 to 4 being “no damage”, “minor damage”, “major damage”, and “destroyed”, respectively.
Overall F1 Score (): This reflects the combined performance of the model in the tasks of building location and damage classification. It is defined as
IV-C Comparison Methods
To show the superiority of our Mamba-based architecture in CD tasks, we select a number of of the state-of-the-art methods based on CNN and Transformer architectures for benchmarking across three distinct tasks: BCD, SCD, and BDA.
For BCD, our comparison methods includes a suite of CNN-based methods: FC-EF [22], FC-Siam-Diff [22], FC-Siam-Conc [22], SiamCRNN [17] (utilizing ResNet architectures ranging from ResNet-18 to ResNet-101 [67]), SNUNet [40], DSIFN [23], HANet [41], and CGNet [42]. For the Transformer-based methods, we compare against ChangeFormer (versions 1 to 6) [30], BIT (with ResNet-18 to ResNet-101 as the backbone for feature extraction) [28], TransUNetCD [33], and SwinSUNet [32].
For the SCD arena, CNN-based competitors encompass HRSCD (variants S1 to S4) [45], ChangeMask [24], SSCD-1 [47], Bi-SRNet [47], and TED [55]. For Transformer-based methods, we evaluate against SMNet [54] and ScanNet [55].
Lastly, for the BDA task, we selected Siamese-UNet [51], MTF [52], and ChangeOS [3] (employing ResNet-18 to ResNet-101) as CNN-based benchmarks. DamFormer [31] stands as the solitary Transformer-based comparison method.
For the above methods, if the detection performance is reported in the original paper on the corresponding dataset, we directly adopt the accuracy in the original paper. If not, we train and test the method on the corresponding dataset based on the hyperparameters recommended in the original paper, using the loss function and data augmentation methods consistent with our method.
IV-D Detection Results and Benchmark Comparison in Three Tasks
@@ IV-D1 BCD
Tables III to V list the performance in BCD of the proposed MambaBCD architecture and some comparison methods on the three benchmark datasets. In these tables, we differentiate between the types of these BCD architectures into CNN-based methods, Tranformer-based methods, and our Mamba-based methods. It can be seen that our approaches significantly outperform either the CNN-based approaches or the Transformer-based architectures. The proposed MambaBCD-Base and MambaBCD-Small architectures achieve the highest and second-highest OA, F1, IoU, and KC for both category-agnostic CD (SYSU dataset) and single-category CD tasks (LEVIR-CD+ and WHU-CD datasets), fully demonstrating the potential of the Mamba architecture for the BCD task. Fig. 7 to Fig. 9 then show some binary change maps predicted by our three architectures on the test sets of the three datasets. It can be seen that the proposed methods accurately detect changes with varying types, scales, and numbers contained in these image pairs, very close to the change reference maps.
@@
| MamabaSCD-Base | 88.12 | 64.03 | 73.68 | 24.11 |
IV-D2 SCD
Table VII compares the SCD performance of the proposed MamabaSCD architectures and representative SCD approaches. Compared to the BCD task, the SCD task requires the change detector capable of effectively modeling not only the spatio-temporal relationships of multi-temporal images, but also learning representative and discriminative representations for accurate land-cover mapping. From Table VII, we can see that the MambaSCD architecture can satisfy both requirements well. The MambaSCD-Small architecture surpasses the Transformer-based SOTA approach, ScanNet, on all four metrics of the SCD task, with OA of 88.38, F1 of 64.10, mIoU of 73.61, and SeK of 24.04 . Note that our methods have achieved the current accuracy without employing loss functions such as those proposed for the SCD task in ScanNet [55].
Based on the SCD results, the detailed “from-to” semantic change information can be obtained, which is presented in Fig. 10. Firstly, we can find that the “from-to” transition matrices obtained by our methods are very close to the ground truth, demonstrating the effectiveness of our MambaSCD architecture again. Then, based on the contents of matrices, we can find that the three semantic changes with the highest percentage between the two temporal phases are “non-vegetated ground surface building”, “low vegetation building”, “low vegetation non-vegetated ground surface”. These results indicate that the areas covered by the SECOND dataset was in the process of urbanization during this period.
@@ IV-D2 SCD
| MamabaSCD-Tiny | 21.51 | 73.42 | |
| MamabaSCD-Small | 54.28 | 146.70 | |
| MamabaSCD-Base | 89.99 | 211.55 |
Table X further lists the number of parameters and GFLOPs of these methods. Compared to the Transformer-based ScanNet method, the GFLOPs of our three architectures are lower. Compared to SMNet, another Transformer-based method, the proposed MamabaSCD-Tiny can achieve higher SeK values with lower number of parameters and GFLOPs. These results demonstrate the effectiveness of the proposed methods and the potential of the Mamba architecture on the SCD task.
@@ IV-D3 BDA
| Siamese-UNet [51] | 85.92 | 65.58 | 71.68 | 86.74 | 50.02 | 64.43 | 71.68 | |
| MTF [52] | 83.60 | 70.02 | 74.10 | 90.60 | 49.30 | 72.20 | 83.70 | |
@@
In order to further show the misclassification between the different damage levels in more detail, we show their confusion matrices in Fig. 12. As can be seen from the confusion matrices, the categories “No damage” and “destroyed” are classified with high accuracy and are not easily misclassified into the remaining two categories. In contrast, the categories of “minor damage” and “major damage” are easily misclassified as other categories, especially the “minor damage” category. A large portion of the “minor damage” category is misclassified as “no damage” and “major damage”. There are mainly the following reasons. Firstly, the optical images in the xBD dataset were acquired with a relatively small viewing angle of the sensor. In many cases, “minor damage” is simply damage to the structure of the house. Little information can be obtained from the orthorectified optical image for such cases. In addition, there is a semantic ambiguity in the definition of the categories minor damage and major damage. We argue that it is difficult to address this problem by starting from the model architecture perspective. It can be considered to introduce of auxiliary information, such as height information or multi-view imagery [21, 69], etc. to help the model make more accurate classification.
| Type | Method | Params | GFLOPs |
| Siamese-UNet [51] | 51.46 | 138.88 | |
| MTF [52] | 44.40 | 268.76 | |
| ChangeOS-18 [3] | 23.49 | 229.22 | |
| ChangeOS-34 [3] | 33.60 | 255.85 | |
| ChangeOS-50 [3] | 52.62 | 469.41 | |
| ChangeOS-101 [3] | 71.61 | 508.41 | |
| DamFormer [31] | 32.52 | 169.34 | |
| MamabaBDA-Tiny | 19.74 | 59.57 | |
| MamabaBDA-Small | 52.11 | 130.80 | |
| MamabaBDA-Base | 87.76 | 195.43 |
Table X lists the the number of parameters and GFLOPs of these BDA approaches. For MambaBDA-Small, our architecture yields more accurate building damage assessment results compared to Siamese-UNet, ChangeOS-101 and MTF, with an improvement in from 5.64 to 9.46 for a similar number of parameters and lower GFLOPs.
@@ IV-E Different Spatio-Temporal Fusion Methods
Modelling spatio-temporal relationships is significant to CD tasks [46, 17, 24]. To demonstrate the effectiveness of our proposed modelling mechanisms, we compare it with several commonly used spatial-temporal modeling methods in CD. They are concatenation operations with feature pyramid network (FPN) [24] (baseline), temporal-symmetric transformation based on 3D convolutional layers (TST) [24], RNN-based method [17], and Tranformer-based method [28]. Table XI lists the performance of these modeling methods on the three CD subtasks. On the simpler BCD tasks, Transformer does not yield significant advantages in performance compared to RNN. However, on the more difficult SCD and BDA tasks, the advantages of the Transformer architecture are illustrated. Compared with these current representative spatio-temporal relationship modeling approaches, our proposed modeling approach based on the Mamba architecture, which inherits the advantages of RNN and Transformer, shows superior performance on the three CD tasks of BCD, SCD, and BDA. Compared to the Transformer-based spatio-temporal modeling approach, our method improves the F1 value by 0.77 on BCD, SeK by 0.6 on the SCD task, and by 1.10 on the BDA task.
We further visualize the features extracted by the MambaBCD architecture in Fig. 14. In the features extracted by the encoder, the features in the region of high response values are not necessarily regions of interest, i.e., changed regions. Whereas, in the decoder, after the spatio-temporal relationships are gradually modeled by the proposed spatio-temporal modelling mechanisms, changed regions gradually exhibit high response values.
@@ IV-F Comparison to Other Backbone Networks
| Encoder | OA | F1 | IoU | KC | Param (M) | GFLOPs |
| ResNet-101 [67] | 91.10 | 81.10 | 68.22 | 75.28 | 46.90 | 131.36 |
| EfficientNet-B5 [70] | 91.60 | 82.26 | 69.86 | 76.84 | 31.63 | 76.88 |
| MixFormer-v3 [35] | 87.70 | 75.00 | 60.00 | 66.87 | 47.03 | 82.98 |
| Swin-Small [34] | 92.09 | 82.22 | 69.80 | 77.16 | 37.75 | 97.99 |
| VMamba-Small [38] | 92.35 | 82.83 | 70.70 | 77.94 | 49.94 | 114.82 |
@@
Finally, we simply compare the Mamba with some representative backbone networks in CD tasks on the SYSU dataset in Table XII. It can be seen that Mamba outperforms the two Transformer backbone networks, Swin-Transformer and MixFormer. This is because, in order to reduce the computational overhead of self-attention operation, these two Transformer backbone networks either adopt a local attention mechanism (Swin-Transformer) [34] or a strided convolutional layer (MixFormer) [35] to reduce the size of the feature maps, which will negatively affect the learning of global contextual information. The Mamba architecture, on the other hand, can reduce its computational consumption to a linear relationship with the number of tokens, thus eliminating the need for these methods, which cause information loss, and allowing the fully learning of global context information. As for the remote sensing change detection task, due to the multi-scale phenomenon of land-cover features and the differences in spatial resolution of different sensors, the sizes of changed objects often vary greatly. Therefore, the features extracted by Mamba through fully learning the global contextual information, are particularly important for the subsequent detection of these changed objects with different sizes.
IV-G GPU Memory Footprint
In Section IV-D, we list the number of parameters and GFLOPs of different models. In this subsection, we further compare the GPU memory footprint of some models. Fig. 15 compares the GPU memory footprint of some CNN-based and Transformer-based methods with our MambaBCD architecture for varying input size. We can see that the trend of GPU memory footprint with input size for MambaBCD-Small is similar to some of the advanced CNN architectures (DSIFN and SNUNet). However, our MambaBCD is able to learn global spatial contextual information as well as adequately model spatio-temporal relationships, thereby exhibiting better detection performance. The GPU memory footprint of MambaBCD-Small is significantly lower compared to ChangeFormerV5, which employs the Transformer architecture as its backbone. Note that ChangeFormerV5 itself has employed strided convolutional layer to reduce the computational overhead of self-attention operations. Also, we can find that the lightweight architecture ChangeFormerV2 has surpassed the Mamba architecture and the two CNN architectures, DSIFN and SNUNet, in terms of the increasing trend of GPU memory footprint when the input size is increased from 1536 to 2048. These results imply that Transformer to be applicable to large scale scenarios of remote sensing images often comes at the cost of information loss. In comparison, Mamba architecture can avoid this problem.
IV-H Robustness Against Degraded Input Data
Finally, we test the robustness of our method to degraded input data. In the inference phase, the data input to the model are subjected to different levels of Gaussian blurring, Gaussian noise, and scaling of the data by different proportions, respectively. Fig. 16 shows the degradation on F1 score of different methods with degraded data on the LEVIR-CD+ dataset. Firstly, we can see that the CNN-based methods are very poorly robust to Gaussian blurred data. F1 scores of SiamCRNN-101 and DSIFN drop by nearly 60 when increasing the standard deviation in the Gaussian blur from 0 to 2. BIT-101 also drops its F1 score by more than 34 due to the adoptation of ResNet to extract features. Whereas the two methods using Transformer as backbone, SwinTransUNet and ChangeFormerV5 are relatively robust to Gaussian blurred data, with F1 scores decreased by 14.83 and 18.43 , respectively. In contrast, the F1 score for our MambaBCD architecture drops by only 8.73.
Fig. 16-(b) shows the effect of adding different levels of Gaussian noise to the input data. With the exception of ChangeFormerV5 and our MambaBCD, the remaining methods have significant F1 decreases ranging from 40.83 to 15.07. Our MambaBCD architecture, on the other hand, shows robustness to Gaussian noise, with only an 8.26 drop in the F1 score for Gaussian noise with a standard deviation of 20.
Fig. 16-(c) compares the robustness of models trained at a single scale to test data at different scales. The F1 value of MambaBCD decreases by only 2.88 when the scaling is increased from 1 to 1.5. When the scaling ratio decreases from 1 to 0.75, the F1 value of MambaBCD decreases by only 1.18. When the scaling factor is reduced to 0.5, all methods, including MambaBCD, show a larger decrease in F1 score. However, compared to the decline of 14.99 of SwinTransUNet and decline of 16.41 of ChangeFormerV5, the decline of MambaBCD decline is smaller, at 12.03 . All the above results again demonstrate the superiority of our proposed Mamba-based CD architecture.